# BATIS: Bayesian Approaches for Targeted Improvement of Species Distribution Models

### v1

@@ 6 Conclusion
SDMs are widely used to inform decisions in conservation, land use, and other areas of impact (Jetz et al. 2019). We hope that BATIS can improve the accuracy of such SDMs in applications where new observations are continually arriving, such as the monitoring of protected areas and endangered species, as well as guiding policies in underresourced or remote regions with lower data availability. Our framework is designed to be readily integrated into existing workflows for creation and utilization of SDMs. In addition, traditional barriers to integration of deep learning into ecology have included computational cost and lack of perceived interpretability. Our proposed approach is computationally lightweight and allow for improved performance without costly retraining, while the focus on uncertainty quantification makes model predictions more transparent, as well as bringing the output closer to the statistical tools already familiar to those working with SDMs.
Limitations.
Effectively combining large-scale ecological patterns with local information remains challenging when the prior variance is very large. In such cases, as shown in Figure 2, the initial predictions of an uncertainty-aware SDM can be overwritten in a first Bayesian update, which can lead to an initial drop in performance before enough updates are performed. Ideally, we would like performance to increase monotonically with number of local observations. Additionally, uncertainty estimation cannot fully correct for the structured biases present in citizen science data, such as observer expertise or the day of the week, which may affect the accuracy of the underlying numbers used as ground truth. Developing more effective strategies to account for these biases remains an active subject of inquiry in statistical ecology (Rocchini et al. 2023). We provide an extended overview of the limitations of our approach in Appendix B.
@@
Future work includes extending BATIS to guide sampling effort in citizen science initiatives, in collaboration with ecologists, and investigating the behavior of different uncertainty estimation approaches with presence-only data (which are often more abundant and readily available than presence-absence checklist data ; Zbinden et al. (2024)). Finally, an impactful extension of our work would be to control for additional sources of biases in the data, e.g. building on Sierra et al. (2025) by considering subdatasets labeled according to different types of bias.
@@ Appendix A Dataset
We used the 19 bioclimatic variables of the WorldClim model Fick and Hijmans (2017). For the USA-Summer and USA-Winter subsets, we followed Teng et al. (2023); Cole et al. (2020) and considered data from WorldClim 1.4, which has data aggregated from 1960-1990. For Kenya and South Africa, we considered data from WorldClim 2.1 instead, which has data aggregated from 1970-2000. Both WorldClim 1.4 and WorldClim 2.1 share the same variables and the same resolution (approximately 1km per pixel). Table 4 summarizes the variables we used.
@@
Table 5 describes the number of hotspots and the number of checklists associated with the four subsets of our dataset, for each split (train/val/test). These splits were obtained following the process described in section 4. Similarly to (Teng et al. 2023) and following recommendations in the literature (Roberts et al. 2017), we pre-processed the data based on the geospatial coordinates of each hotspot in order to prevent auto-correlation and over-fitting using DBSCAN to cluster hotspots within 5km of each other and ensuring that individual members of each cluster were proportionally distributed across splits.
@@
We limited the species list to the 1054 species regularly found in Kenya according to Avibase (Avibase - The World Bird Database ).
@@
We limited the species list to the 755 species regularly found in South Africa according to Bird Life South Africa (BirdLife - Bird Checklists of South Africa ).
@@
We limited the species list to the 670 species regularly found in continental United States according to the 2022 Checklist of the Continental United States from the American Birding Association (American Birding Association ; Checklists of the continental United States ). More specifically, we only considered species denoted by ABA Codes 1 or 2, which include regular breeding species and visitors that are widespread across the country. We did not consider the species that are known to only occur in Alaska, Hawaii or the U.S. Territories (e.g. Guam, Puerto Rico, American Samoa, etc.).
@@
We provide a visual overview of the Kenya subdataset below. Figure 4 shows the geographical distribution of hotspots, Figure 5 shows the geographical distribution of species, and 6 shows the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the USA-Winter subdataset below. Figure 8 shows the geographical distribution of hotspots, Figure 9 shows the geographical distribution of species, and 7 shows the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the South-Africa subdataset below. Figure 10 shows the geographical distribution of hotspots, Figure 11 shows the geographical distribution of species, and 12 shows the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the USA-Winter subdataset below. Figure 13 shows the geographical distribution of hotspots, Figure 14 shows the geographical distribution of species, and 15 shows the distribution of non-zero encounter rates per hotspot.
Figure 4: Geographical distribution of the total number of hotspots associated with our Kenya subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the total number of hotspots located within that cell.
@@
Figure 5: Geographical distribution of the total number of non-zero encounter rates associated with our Kenya subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the mean of non-zero encounter rates across hotspots located within that cell.
Figure 6: Distribution of the number of species encountered per hotspot for the Kenya subset, for a) Training and b) Test.
Figure 7: Distribution of the number of species encountered per hotspot for the USA-Winter subset, for a) Training and b) Test.
Figure 8: Geographical distribution of the total number of hotspots associated with our USA-Winter subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the total number of hotspots located within that cell.
@@
Figure 9: Geographical distribution of the total number of non-zero encounter rates associated with our USA-Winter subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the mean of non-zero encounter rates across hotspots located within that cell.
Figure 10: Geographical distribution of the total number of hotspots associated with our South Africa subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the total number of hotspots located within that cell.
@@
Figure 11: Geographical distribution of the total number of non-zero encounter rates associated with our South Africa subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the mean of non-zero encounter rates across hotspots located within that cell.
Figure 12: Distribution of the number of species encountered per hotspot for the South Africa subset, for a) Training and b) Test.
Figure 13: Geographical distribution of the total number of hotspots associated with our USA-Summer subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the total number of hotspots located within that cell.
@@
Figure 14: Geographical distribution of the total number of non-zero encounter rates associated with our USA-Summer subdataset for a) Training and b) Test. The value assigned to each 40km2 grid cell corresponds to the mean of non-zero encounter rates across hotspots located within that cell.
Figure 15: Distribution of the number of species encountered per hotspot for the USA-Summer subset, for a) Training and b) Test.
Appendix B Limitations
@@
There remains tension between the broad environmental variables learned by deep SDMs (our Bayesian prior) and the hotspot-specific features captured by observational information (our updates to that prior). As shown in Figure 2, the initial predictions of an uncertainty-aware SDM can be overwritten in a first Bayesian update, when the estimated prior variance is large. This can lead to an initial drop in performance before enough updates are performed to boost performance, even though performance would ideally increase monotonically with number of updates. A simple approach to mitigate such behavior is to consider a weighted average of predictions from the prior and the updated SDM, progressively giving more importance to the updates with more observations.
Citizen science data is heavily influenced by predictable factors such as the day of the week, the time of day, the expertise of the observer, the weather, and the proximity to population centers (Sierra et al. 2025). Such biases are structured, which technically violates the randomness assumption of aleatoric uncertainty. Epistemic uncertainty estimation approaches can potentially correct for biases if they can be explicitly modeled, but are unable to do so if they are not encoded in the data. While our work shows that is possible to compensate partially for data gaps associated with the irregular distribution of citizen science data, some biases may affect the accuracy of the underlying numbers used as ground truth, which cannot be compensated for by such methods and remain an active object of inquiry in statistical ecology (Rocchini et al. 2023).
Appendix C Methods
@@
A Multi-Layer Perceptron (MLP) is a simple feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer (Goodfellow, Bengio, and Courville 2016). We used a single hidden layer of dimension 64 and sigmoid activation layers. The MLP was trained with the binary cross-entropy loss, i.e.
@@
We used a conventional Resnet-18 architecture (He et al. 2016), initialized with weights pre-trained on ImageNet. The model was trained with the binary cross-entropy loss (equation 6), and implemented in Pytorch with the help of the torchvision.models.resnet18 class.
@@
Monte-Carlo Dropout (MCD) randomly deactivates a proportion $\rho$ of neurons at training and test time, and uses $M$ forward passes to compute the mean and variance for each hotspot. Dropout layers need to be careful placed, as they can quickly degrade accuracy for large models, which will dramatically increase the computation time required to achieve a satisfying performance (Kendall, Badrinarayanan, and Cipolla 2016). We did the same as (Kendall and Gal 2017) and added only a single dropout layer before the last layer of the network.
@@
In order to optimize performance, we followed recommendations formulated in (Sluijterman, Cator, and Heskes 2023) and added a separate regularization of the mean and the variance estimate. Our loss function therefore became
@@
Heteroscedastic Regression Neural Network (Kendall and Gal 2017).
@@ Appendix D Experiments
All deep learning models were trained using the Adam (Kingma 2014) optimizer, with a batch size of 128 and for a maximum of 50 epochs. We trained each model with three different seeds, and we used 30 dropout pass at test time to compute the mean and variance of predictions for the approaches requiring a dropout layer to quantify uncertainty (Gal and Ghahramani 2016). Random noise, blur and vertical/horizontal flipping were used for data augmentation, and we normalized inputs (satellite imagery and/or bioclimatic variables) using training set statistics. The optimizer hyperparameter values we used in each of our experiments are listed in Table 8. In addition, a summary of the hyperparameter values we used for each of our uncertainty-aware methods can be found in Table 7.
We report results generated with the same learning rate ($0.0001$) for all Resnet18-based models because it was demonstrated to be the one that gave the best results in the related SatBird task (Teng et al. 2023). For our Monte-Carlo Dropout approach, we took inspiration from the experiments in Gal and Ghahramani (2016) and considered $\rho=0.2$ as the dropout rate for MCD and HetReg. For the warmup, $\lambda_{\mu}$ and $\lambda_{\sigma}$ hyperparameters of our Mean-Variance Network and HetReg approaches, we considered $\text{{warmup}}=\left[2,5,10\right]$, $\lambda_{\mu}=\left[0.1,0.001,0.0001\right]$, $\lambda_{\sigma}=\left[0.1,0.001,0.0001\right]$, and found that using $\text{{warmup}}=5$, $\lambda_{\mu}=0.1$ and $\lambda_{\sigma}=0.1$ led to the best results.
@@
Our code was developed in Python 3.10. The required packages can be installed using the requirements.txt file provided in our Github repository. We validated that all our code can be executed on :
  • SLURM-based clusters
  • macOS Sonoma 14.1 (with Apple M Series chips)
  • Windows 11 (with NVIDIA GeForce RTX 4070)
Execution environment
Model Training (Resnet-18).
Model Training (Baselines).
Bayesian Updating Framework.
@@
Appendix E Results
@@
Appendix F HuggingFace Repository
Appendix G Code Repository
Appendix H Licenses
@@
Appendix I Reproducibility Checklist
We provide below further explanations for some of our reproducibility checklist answers.
Question 3.3 : All novel datasets introduced in this paper are included in a data appendix.
Due to file size limitations, we only included a subset of the dataset in the data appendix. The complete dataset is available in the anonymous HuggingFace repository (see page 1).
Question 4.2 : This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting.
Yes, in the technical appendix.
Question 4.3 : Any code required for pre-processing data is included in the appendix.
Yes. It is also available on the provided Github link (see page 1).
Question 4.4 : All source code required for conducting and analyzing the experiments is included in a code appendix.
Yes. It is also available on the provided Github link (see page 1).
Question 4.7 : If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results.
Yes. It is briefly described in the main paper, and we provide more details in the technical appendix.
Question 4.8 : This paper specifies the computing infrastructure used for running experiments (hardware and soft-ware), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks.
Yes. Details are provided in the technical appendix.
Question 4.13 : This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments.
Yes. Details are provided in the technical appendix.

### v2

@@ 7 Conclusion
SDMs are widely used to inform decisions in conservation, land use, and other areas of impact (sdmessentialtool). We hope that BATIS can improve the accuracy of such SDMs in applications where new observations are continually arriving, such as the monitoring of protected areas and endangered species, as well as guiding policies in underresourced or remote regions with lower data availability. Our framework is designed to be readily integrated into existing workflows for creation and utilization of SDMs. In addition, traditional barriers to integration of deep learning into ecology have included computational cost and lack of perceived interpretability. Our proposed approach is computationally lightweight and allow for improved performance without costly retraining, while the focus on uncertainty quantification makes model predictions more transparent, as well as bringing the output closer to the statistical tools already familiar to those working with SDMs.
@@
Future work includes extending BATIS to guide sampling effort in citizen science initiatives, in collaboration with ecologists, and investigating the behavior of different uncertainty estimation approaches with presence-only data (which are often more abundant and readily available than presence-absence checklist data ; zbindenpresenceabsence). Finally, an impactful extension of our work would be to control for additional sources of biases in the data, e.g. building on sierra2025divshift by considering subdatasets labeled according to different types of bias.
@@ Appendix A Dataset
We used the 19 bioclimatic variables of the WorldClim model worldclim. For the USA-Summer and USA-Winter subsets, we followed NEURIPS2023_ef7653bb; geolifeclef and considered data from WorldClim 1.4, which has data aggregated from 1960-1990. For Kenya and South Africa, we considered data from WorldClim 2.1 instead, which has data aggregated from 1970-2000. Both WorldClim 1.4 and WorldClim 2.1 share the same variables and the same resolution (approximately 1km per pixel). Table 4 summarizes the variables we used.
@@
Table 5 describes the number of hotspots and the number of checklists associated with the four subsets of our dataset, for each split (train/val/test). These splits were obtained following the process described in section 4. Similarly to (NEURIPS2023_ef7653bb) and following recommendations in the literature (autocorrgis), we pre-processed the data based on the geospatial coordinates of each hotspot in order to prevent auto-correlation and over-fitting using DBSCAN to cluster hotspots within 5km of each other and ensuring that individual members of each cluster were proportionally distributed across splits.
@@
We limited the species list to the 1054 species regularly found in Kenya according to Avibase (Avibase).
@@
We limited the species list to the 755 species regularly found in South Africa according to Bird Life South Africa (BirdLife).
@@
We limited the species list to the 670 species regularly found in continental United States according to the 2022 Checklist of the Continental United States from the American Birding Association (ABA). More specifically, we only considered species denoted by ABA Codes 1 or 2, which include regular breeding species and visitors that are widespread across the country. We did not consider the species that are known to only occur in Alaska, Hawaii or the U.S. Territories (e.g. Guam, Puerto Rico, American Samoa, etc.).
@@
We provide a visual overview of the Kenya subdataset below. Figure 4 shows the geographical distribution of hotspots and species, as well as the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the USA-Winter subdataset below. Figure 5 shows the geographical distribution of hotspots and species, as well as the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the South-Africa subdataset below. Figure 6 shows the geographical distribution of hotspots and species, as well as the distribution of non-zero encounter rates per hotspot.
@@
We provide a visual overview of the USA-Winter subdataset below. Figure 7 shows the geographical distribution of hotspots and species, as well as the distribution of non-zero encounter rates per hotspot.
Appendix B HuggingFace Repository
Appendix C Code Repository
Appendix D Licenses
@@
Figure 4: From left to right, for a) Training and b) Test : Geographical distribution of the total number of hotspots (left) and non-zero encounter rates (middle) associated with our Kenya subdataset, and distribution of the number of species encountered per hotspot (right).
@@
Figure 5: From top to bottom, for a) Training and b) Test : Geographical distribution of the total number of hotspots (top) and non-zero encounter rates (middle) associated with our US-Winter subdataset, and distribution of the number of species encountered per hotspot (bottom).
@@
Figure 6: From left to right, for a) Training and b) Test : Geographical distribution of the total number of hotspots (left) and non-zero encounter rates (middle) associated with our South-Africa subdataset, and distribution of the number of species encountered per hotspot (right).
@@
Figure 7: From top to bottom, for a) Training and b) Test : Geographical distribution of the total number of hotspots (top) and non-zero encounter rates (middle) associated with our US-Sumer subdataset, and distribution of the number of species encountered per hotspot (bottom).
Figure 8: Iterative improvements of MAE, MSE and Top-10 metrics on the Kenya sub-dataset with increasing number of BATIS updates for our best performing uncertainty-aware model (HetReg), with and without progressive blending of prior and posterior predictions. Results are shown for various values for hyperparameter $\lambda$ (see equation 6).
Appendix E Limitations
@@
As shown in Figure 2, the initial predictions of an uncertainty-aware SDM can be overwritten in a first Bayesian update, when the estimated prior variance is large. This can lead to an initial drop in performance before enough updates are performed to boost performance, even though performance would ideally increase monotonically with number of updates. A simple approach to mitigate such behavior is to consider a weighted average of predictions from the prior and the updated SDM, progressively giving more importance to the updates with more observations. For example, we can consider blended predictions
$p_{blend}=(1-w_{t})\cdot p_{prior}+w_{t}\cdot p_{posterior}$ (6)
where $w_{t}\in(0,1]$ is an appropriate discounting factor. A practical choice for $w_{t}$ is $w_{t}=1-\exp^{-\lambda t}$, where $t$ corresponds to the number of BATIS posterior updates and $\lambda\in(0,1]$ is a tunable hyperparameter. Figure 8 illustrates how this strategy can progressively smooth performance curves for our best performing uncertainty-aware model (HetReg) on the Kenya sub-dataset.
Appendix F Methods
@@
A Multi-Layer Perceptron (MLP) is a simple feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer (deeplearningbook). We used a single hidden layer of dimension 64 and sigmoid activation layers. The MLP was trained with the binary cross-entropy loss, i.e.
@@
We used a conventional Resnet-18 architecture (resnetcvpr), initialized with weights pre-trained on ImageNet. The model was trained with the binary cross-entropy loss (equation 7), and implemented in Pytorch with the help of the torchvision.models.resnet18 class.
@@
Monte-Carlo Dropout (MCD) randomly deactivates a proportion $\rho$ of neurons at training and test time, and uses $M$ forward passes to compute the mean and variance for each hotspot. Dropout layers need to be careful placed, as they can quickly degrade accuracy for large models, which will dramatically increase the computation time required to achieve a satisfying performance (manual2). We did the same as (hetregpaper) and added only a single dropout layer before the last layer of the network.
@@
In order to optimize performance, we followed recommendations formulated in (mvapaper1) and added a separate regularization of the mean and the variance estimate. Our loss function therefore became
@@
Heteroscedastic Regression Neural Network (hetregpaper).
@@ Appendix G Experiments
All deep learning models were trained using the Adam (kingma2014adam) optimizer, with a batch size of 128 and for a maximum of 50 epochs. We trained each model with three different seeds, and we used 30 dropout pass at test time to compute the mean and variance of predictions for the approaches requiring a dropout layer to quantify uncertainty (pmlr-v48-gal16). Random noise, blur and vertical/horizontal flipping were used for data augmentation, and we normalized inputs (satellite imagery and/or bioclimatic variables) using training set statistics. The optimizer hyperparameter values we used in each of our experiments are listed in Table 8. In addition, a summary of the hyperparameter values we used for each of our uncertainty-aware methods can be found in Table 7.
We report results generated with the same learning rate ($0.0001$) for all Resnet18-based models because it was demonstrated to be the one that gave the best results in the related SatBird task (NEURIPS2023_ef7653bb). For our Monte-Carlo Dropout approach, we took inspiration from the experiments in pmlr-v48-gal16 and considered $\rho=0.2$ as the dropout rate for MCD and HetReg. For the warmup, $\lambda_{\mu}$ and $\lambda_{\sigma}$ hyperparameters of our Mean-Variance Network and HetReg approaches, we considered $\text{{warmup}}=\left[2,5,10\right]$, $\lambda_{\mu}=\left[0.1,0.001,0.0001\right]$, $\lambda_{\sigma}=\left[0.1,0.001,0.0001\right]$, and found that using $\text{{warmup}}=5$, $\lambda_{\mu}=0.1$ and $\lambda_{\sigma}=0.1$ led to the best results.
@@
Our code was developed in Python 3.10. The required packages can be installed using the requirements.txt file provided in our Github repository.
@@
Execution Environment
We validated that all our code can be executed on :
  • SLURM-based clusters
  • macOS Sonoma 14.1 (with Apple M Series chips)
  • Windows 11 (with NVIDIA GeForce RTX 4070)
Model Training (Resnet-18).
Model Training (Baselines).
Bayesian Updating Framework.
Appendix H Results
@@
Figure 11: Iterative improvements for the different uncertainty estimation approaches with increasing number of checklist updates for the MAE, MSE and Top-10 metrics on the USA-Winter Region test set. We report the mean on three seeds and standard deviations for each model.